OpenStudy (anonymous):

Find the remaining factors for x4 - 10x3 + 35x2 - 50x + 24 if two factors are x - 1, and x - 3 Polynomial Functions.

OpenStudy (triciaal):

@ziqbal103

OpenStudy (campbell_st):

well this is a quick and dirty solution... and involves equating coefficients. \[(x -1)(x -3_ = x^2 - 4x + 3 \] then the other quadratic factor is \[(x + a)(x + b) = x^2 + (a + b)x + ab\] then product of the 2 quadratic is = x^4 + (-4 + a + b)x^3 +(3 +ab -4(a+b))x + (-4ab+3(a+b))x + 3ab\] comparing coeffiecients constant 24 = 3ab so ab = 8 the coefficient of x^3 -10 = -4 + a + b a + b = -6 you can check the values of ab and a + b by substituting into the coefficients of x and x^2 so the 2nd quadratic is \[x^2 -6x + 8\] factor this for the other binomial factors.

OpenStudy (anonymous):

\[x^4-10 x^3+35 x^2-50 x+24=(x-4) (x-3) (x-2) (x-1) \]